A particle perform simple harmonic motion in a straight line about a center C in the line.When it's distances from C are 5.0 and 13.0cm,it's velocities are 25.0 and 7.0cm/s respectively.Find the period and the amplitude of the oscillation. Hint: use V=±w(angular velocity)√(A² - y²) and solve simultaneously.?

1
Expert's answer
2022-03-08T08:55:32-0500

"v_1=\omega\sqrt{A_0^2-A_1^2},\\
v_2=\omega\sqrt{A_0^2-A_2^2},"

"\omega=2,"

"A=13.5~cm,"

"T=\frac{2\pi}{\omega}=3.14~s."